# Shadow of black holes with consistent thermodynamics

### v1

@@ IV.1 Shadow critical curves
It turns out that the geodesic equations for the spacetime (1) with a running $G(r,M,a)$ are separable because of the existence of a Carter-like constant. This can be demonstrated using the Hamilton-Jacobi approach. Let us consider the following Lagrangian
@@
where the dot represents the derivative with respect to the affine parameter $\lambda$. Due to the axisymmetry and stationarity of the spacetime (1), there are two conserved quantities along the geodesics $E=-p_{t}$ and $L=p_{\varphi}$, which correspond to the energy and the azimuthal angular momentum on the geodesic. The conjugate momenta $p_{\mu}\equiv\partial\mathcal{L}/\partial\dot{x}^{\mu}$ are defined in the standard way. One can then invert $E$ and $L$ to get the expressions for $\dot{t}$ and $\dot{\varphi}$:
@@
where $\mathcal{S}$ is the Jacobi action. It turns out that the metric (1) allows for a separable $\mathcal{S}$ as a solution to Eq. (24):
@@
$\frac{\left[\left(r^{2}+a^{2}\right)E-aL\right]^{2}}{\Delta}+\left(L-aE\right)^{2}+\Delta\left(\frac{d\mathcal{S}_{r}}{dr}\right)^{2}=\cos^{2}\theta\left(a^{2}E^{2}-\frac{L^{2}}{\sin^{2}\theta}\right)-\left(\frac{d\mathcal{S}_{\theta}}{d\theta}\right)^{2}\,.$ (26)
@@
We thus find that the constants of motion $\xi$ and $\eta$ on a spherical photon orbit can be labeled by the radius $r_{p}$ of the orbit using Eqs. (31) and (32). The orbits that contribute to the critical curve on the image plane should be unstable, i.e., $d^{2}\mathcal{R}(r)/dr^{2}|_{r_{p}}>0$. By using Eqs. (31) and (32), we obtain the identify
@@
In order to visualize the shadow critical curve, we define the celestial coordinates $(\alpha,\beta)$ on the image plane. The coordinate $\alpha$ is the apparent perpendicular distance between the critical curve and the axis of symmetry, i.e., the spin axis. The coordinate $\beta$, on the other hand, is defined by the apparent perpendicular distance between the critical curve and the plane normal to the spin axis, with the axis piercing through the plane at the black hole. We denote the position of the observer by $(r_{0},\theta_{0})$ with $\theta_{0}$ being the inclination angle between the spin axis of the black hole and the direction to the observer. For an observer in the far exterior and asymptotically flat region $r_{0}\rightarrow\infty$, the celestial coordinates $(\alpha,\beta)$ can be expressed as Vazquez:2003zm
$\alpha=\lim_{r_{0}\rightarrow\infty}\left(-r_{0}^{2}\sin\theta_{0}\frac{d\psi}{dr}\Bigg|_{r_{0},\theta_{0}}\right)\,,\qquad\beta=\lim_{r_{0}\rightarrow\infty}\left(r_{0}^{2}\frac{d\theta}{dr}\Bigg|_{r_{0},\theta_{0}}\right)\,.$ (34)
Using the geodesic equations Eqs. (22) and (27), one can rewrite Eq. (34) in terms of the constants of motion $(\xi,\eta)$ and the inclination angle $\theta_{0}$ in the following expressions: [^3] The expressions (35) can also be obtained using the tetrad formalism 1973ApJ…183..237C , by which one can define a generalized version of the celestial coordinates for arbitrary observers Frost:2024gza .
@@ IV.2 Shadow observables
where one integrates over the entire $\alpha$-axis on the image. One can recast this integration in terms of $r_{p}$. This integral then encompasses the entire range of unstable spherical photon orbits that contribute to the critical curve.
@@ V Results
After introducing the derivation of shadow critical curves and the associated observables, we will discuss the shadows of the black holes that satisfy the consistency of black hole thermodynamics in this section. In particular, we will focus on the black hole metric (1) with the running Newton coupling $G(r,M,a)$ parameterized by $G(z)$ with $z$ given in Eq. (16). We will consider the following sets of models: (l,p)=(1,1)(l,p)=(1,-1), $(1,0)$, $(1,4)$, and $(0,1)$. The main goal is twofold: one is to place constraints on the models using the observations of black hole images, and the other is to look for possible features in the shadow contours that are shared among these models.
@@ V.1 Corrections without mass suppression (l,p)=(1,1)(l,p)=(1,-1)
where the indices $(m,n)\in\mathbb{Z}$. In this parameterization (42), we have a set of coupling constants $(a_{i},b_{i})$ which do not depend on $M$ and are truly dimensionless. As we mentioned, the constraining power on $(a_{i},b_{i})$ will not be suppressed by the black hole mass $M$ when considering supermassive black hole images.
@@
Figure 1: The shadow critical curves (Left) for $b_{1}=a_{2}=0$ and different values of $b_{2}$, and (Right) for $b_{1}=b_{2}=0$ and different values of $a_{2}$.
Figure 2: (Left) The critical curves for fixed $a_{2}=0$; the blue and the red curves with $(b_{1},b_{2})=(-1,0.1)$ and $(b_{1},b_{2})=(-1,-1)$, respectively. (Right) The critical curves for fixed $b_{2}=0$; the blue and the red curves with $(b_{1},a_{2})=(-1,-0.1)$ and $(b_{1},a_{2})=(-1,1)$, respectively. In each panel, the shaded regions represent the results of contours scanned when taking $b_{1}\geq-1$. For $b_{1}\rightarrow\infty$, the critical curves converge to the Kerr results (black curves).
@@
As one can see from Eq. (42), the choice of the coefficients $(a_{i},b_{i})$ may increase or decrease the Newton coupling $G(z)$, which can directly alter the size of the critical curves. In Fig. 3, we show the fractional diameter deviation $\delta$ in the parameter space $(a_{*},b_{2})$ and $(b_{1},b_{2})$ on the left and right panels, respectively, for $a_{1}=b_{1}$ and $a_{2}=0$. On the left panel, we choose $a_{1}=b_{1}=0$, while on the right panel, we fix $a_{*}=0.9$. The black and green curves indicate the constraints on $\delta$ from the upper limit of the Keck and VLTI measurements, respectively. The parameter space toward the blue region is more preferred observationally.
@@ V.2 Mass-independent scale identification $(l,p)=(1,0)$
Here we consider the model with $(l,p)=(1,0)$, which belongs to the class of mass-independent scale identification in the FRG method. In this case, Eqs. (16) and (19) can be rewritten as
@@
Again, for generality, we assume $G(z)$ to be of the form of Padé expansion with respect to $z$
@@
where the indices $(m,n)\in\mathbb{Z}$ and we have a corresponding set of coupling constants $(c_{i},d_{i})$. [^4] One simple mass-independent scale identification discussed in Ref. Chen:2022xjk corresponds to $c_{n}=0$ for all $n$, $d_{1}=\tilde{\omega}$, and $d_{m}=0$ for $m\neq 1$. The ansatz (46) for $G(z)$ ensures that it recovers Newton constant $G_{0}$ when $|r|\rightarrow\infty$. In fact, when incorporating Eq. (46) into the metric (1), the ADM mass is still given by $M$, and the spacetime satisfies the Solar system constraints Williams:2004qba . Also, the coupling $G(z)$ is symmetric with respect to $r=0$, although the ring on the surface $r=0$ is singular.
@@
Assuming $a_{*}=0.9$ and an edge-on inclination, i.e., $\theta_{0}=\pi/2$, we show the critical curves with different choices of $(c_{1}/G_{0}M^{2},d_{1}/G_{0}M^{2},c_{2}/G_{0}^{2}M^{4},d_{2}/G_{0}^{2}M^{4})$ in Figs. 5 and 6. We first notice that the contributions from $c_{2}$, $d_{2}$, and higher-order coupling constants are subdominant as compared with those from $c_{1}$ and $d_{1}$. One reason for this is because of the mass suppression of these higher-order corrections. Another reason is that the factor $1/(r_{*}^{2}+a_{*}^{2})$ evaluated on the spherical photon orbits is generically smaller than unity, [^5] For example, it is $1/9$ for Schwarzschild black holes, and $1/2$ for extreme Kerr black holes. such that it also suppresses the higher-order corrections. Therefore, from now on, we will focus on the corrections induced by the parameters $c_{1}$ and $d_{1}$.
In addition to the hierarchy between different orders of coefficients, one can see from Figs. 5 and 6 that if the coefficients are chosen in a way such that $G(z)<G_{0}$, the critical curves shrink, acquire more distortion, and display a more D-shaped contour, e.g., the red contour in Fig. 5 where $c_{1}<0$ and the blue contour in Fig. 6 where $d_{1}>0$. This is because, when fixing $a_{*}$, the decrease of the running Newton coupling implies a smaller event horizon, and getting closer to the extreme configuration.
@@ V.4 Spin-independent scale identification $(l,p)=(0,1)$
Finally, we consider the spin-independent scale identification such that the Newton coupling $G(z)$ does not depend on the spin of the black hole. In order to satisfy the consistency of black hole thermodynamics, i.e., Eq. (15), we have $G(r,M,a)=G(Mr)$. Without loss of generality, we choose $(l,p)=(0,1)$ such that
@@
The Newton function is again parameterized using Padé expansion
@@
In Fig. 12, we show the logarithm of $\Delta C$ in the same region of the parameter space as in Fig. 11. Similar to the case of $(l,p)=(1,4)$, in some regions of the parameter space, there is only one root for $\Delta(r)=0$, hence only one horizon exists. These regions correspond to the region above (below) the red curves in the left (right) panel. One can see that the extremality deviation $\epsilon$ can nicely capture the behavior of $\Delta C$, even within the parameter space where there is only one event horizon.
@@ V.5 Hindrance in shrinking shadow size further
In order to demonstrate this in our setup, which has a huge parameter space, we consider different sets of $(l,p)$, then take the spin $a_{*}$ and one of the coupling coefficients in $G(z)$ as two varying parameters, with other coupling coefficients set to zero. For a given $a_{*}$, we tune the value of the varying coupling coefficient to identify the smallest fractional diameter deviation $\delta_{\textrm{min}}$, assuming a D-shaped shadow contour. Even for a given $(l,p)$, when $a_{*}$ changes, the varying coupling coefficient that corresponds to $\delta_{\textrm{min}}$ also varies. In Fig. 13, we demonstrate the smallest fractional diameter deviation $\delta_{\textrm{min}}$ of a D-shaped shadow in terms of $a_{*}$. Each curve represents models with $\theta_{0}=\pi/2$ and different $(l,p)$ as well as the main varying coupling parameter indicated in the legend. Other coupling parameters are set to zero. In each model with a given $a_{*}$, the minimum shadow size occurs when the main varying coupling parameter is tuned such that the black hole spacetime is near the extremal configuration. When $a_{*}=1$, all models share the same $\delta_{\textrm{min}}$ with all the coupling parameters vanishing. This corresponds to the extremal configuration of the Kerr black hole. On the other hand, when $a_{*}=0$, the model of (l,p)=(1,1)(l,p)=(1,-1) with varying $a_{2}$ (magenta solid curve), the model of $(l,p)=(1,0)$ with varying $c_{1}$ (blue solid curve), and the model of $(l,p)=(0,1)$ with varying $g_{2}$ (black solid curve) degenerate because they both have $G(z)/G_{0}-1\propto 1/r^{2}$. Similarly, the model of $(l,p)=(1,0)$ with varying $d_{1}$ (blue dashed curve), the model of $(l,p)=(0,1)$ with varying $h_{2}$ (black dashed curve), and the model of (l,p)=(1,1)(l,p)=(1,-1) with varying $b_{2}$ (magenta dashed curve) degenerate because they all have $G_{0}/G(z)-1\propto 1/r^{2}$. The shaded regions surrounding the curves represent the results scanned for varied $\theta_{0}$.
@@ VI Spinning black holes beyond Kerr bound
In Subsecs. V.3 and V.4 in which we considered the models with $(l,p)=(1,4)$ and $(l,p)=(0,1)$, respectively, we pointed out that there are regions of parameter space where $\Delta(r)=0$ only has one root, hence only one horizon exists. One natural question is: in these cases, can one keep increasing the spin parameter $a_{*}$, go beyond the usual Kerr bound, i.e., $a_{*}\leq 1$, and the event horizon still exists?
@@
It turns out that it is possible to go beyond the Kerr bound, but the shadow images may develop irregular structures. More explicitly, we find that when $a_{*}$ is sufficiently large, a branch of stable spherical photon orbits may appear outside the event horizon. This branch of stable spherical photon orbits connects two distinct branches of unstable photon orbits. In this case, the shadow critical curve develops a cuspy structure, as shown in Fig. 14, i.e., the contours with $a_{*}=3$ and $a_{*}=4$. This can be understood from Fig. 15, where we show the constants of motion $\xi(r_{p})$ and $\eta(r_{p})$ given in Eqs. (31) and (32), respectively. As we discussed in Sec. IV, these constants of motion are functions of the radii $r_{p}$ of different spherical photon orbits. Note that in this figure, the horizontal axis is expressed by the rescaled radius $r_{p*}\equiv r_{p}/G_{0}M$. On the left panel of Fig. 15, we choose the parameters that correspond to the darker green contour in Fig. 14, i.e., $a_{*}=2$. In this case, $\xi(r_{p})$ is a monotonically decreasing function in $r_{p}$, and the orbits that contribute to the contour are all unstable. The shadow critical curve is a smooth and closed contour. On the other hand, on the right panel of Fig. 15, in which the spin $a_{*}=4$ is sufficiently large, $\xi(r_{p})$ has one local minimum and one local maximum, which are connected by a branch of stable spherical photon orbits (the dashed green segment. See also Eq. (33)). This stable branch of orbits connects to two separated branches of unstable orbits on its two ends. There are two distinct unstable orbits, each belonging to the two sides of unstable branches, that contribute to the same point $(\alpha,\beta)$ on the image plane, i.e., the cusps on the contours. Such a discontinuity gives rise to the cuspy structure in the shadow critical curve, and it is a manifestation of the rich spacetime structure in this model, albeit the Liouville integrability of the geodesic dynamics ensured by the Carter constant $\mathcal{K}$ (see Eqs. (27) and (28)). Black hole models with similar cuspy structures in the shadow critical curves can also be found in Refs. Cunha:2017eoe ; Wang:2017hjl . [^8] Cuspy structure in shadow boundary has also been reported in the models in which two distinct branches of unstable photon orbits transit discontinuously, without resorting to stable spherical photon orbits Qian:2021qow . The model considered in Ref. Wang:2017hjl also has Liouville integrable geodesics. The metric line element of the model considered in Ref. Wang:2017hjl can be expressed in the form of Eq. (1) with $G(r)=G_{0}+\bar{\eta}/2Mr^{2}$ and $\bar{\eta}$ a non-Kerr parameter. This Newton coupling satisfies thermodynamic consistency only when $\bar{\eta}\propto 1/M$. If this choice is made, this falls into the class of solutions discussed in Subsec. V.4.
@@
Before closing this section, we would like to comment on the physical implications of this type of cuspy shadows and the corresponding black hole spacetimes. First of all, the shadow contours with such a ``circular sector" shape are very different from the usual D-shaped shadows, and can be easily distinguished through observations. Furthermore, from theoretical points of view, the existence of stable spherical photon orbits implies the possibility of accumulating photons (or gravitons) around the black hole for a long time. Such an accumulation may hint toward some sorts of instability of the entire geometry. The relevant questions are whether such systems are truly unstable, and if so, how long the instability timescale is. However, the answer to these questions requires a thorough numerical analysis that takes into account the nonlinear effects of the underlying gravitational theory, as well as the back reactions induced by the accumulating particles. Such an analysis is beyond the scope of this paper.
@@
Figure 15: (Left) For the black holes whose shadow critical curves do not have cuspy structures, the constant of motion $\xi(r_{p})$ is a monotonic function of $r_{p}$. All the spherical photon orbits that contribute to the shadow critical curve are unstable. (Right) The cuspy structure on the image plane in Fig. 14 is due to the existence of the local maximum and the minimum in $\xi(r_{p})$. In particular, the spherical photon orbits with radii between the two local extrema, i.e., the green dashed line segment, are stable orbits with $d^{2}\mathcal{R}(r)/dr^{2}|_{r_{p}}<0$ (see Eq. (33)). The impact parameters of the orbits in this line segment are represented by the dashed segments in Fig. 14. Note that for $\theta_{0}=\pi/2$, the spherical photon orbits that contribute to the critical curve are those with radii bounded by the two roots of $\eta(r_{p})=0$. The left edge of each panel represents the location of the event horizon, which is very close to the smaller root of $\eta(r_{p})=0$.

### v2

@@ IV.1 Shadow critical curves
It turns out that the geodesic equations for the spacetime (II) with a running $G(r,M,a)$ are separable because of the existence of a Carter-like constant. This can be demonstrated using the Hamilton-Jacobi approach. Let us consider the Lagrangian
@@
where the dot represents the derivative with respect to the affine parameter $\lambda$. Because of the axisymmetry and stationarity of the spacetime (II), there are two conserved quantities along the geodesics $E=-p_{t}$ and $L=p_{\varphi}$, which correspond to the energy and the azimuthal angular momentum on the geodesic. The conjugate momenta $p_{\mu}\equiv\partial\mathcal{L}/\partial\dot{x}^{\mu}$ are defined in the standard way. One can then invert $E$ and $L$ to get the expressions for $\dot{t}$ and $\dot{\varphi}$:
@@
where $\mathcal{S}$ is the Jacobi action. It turns out that the metric (II) allows for a separable $\mathcal{S}$ as a solution to Eq. (25):
@@
$-\frac{\left[\left(r^{2}+a^{2}\right)E-aL\right]^{2}}{\Delta}+\left(L-aE\right)^{2}+\Delta\left(\frac{d\mathcal{S}_{r}}{dr}\right)^{2}$
$=\cos^{2}\theta\left(a^{2}E^{2}-\frac{L^{2}}{\sin^{2}\theta}\right)-\left(\frac{d\mathcal{S}_{\theta}}{d\theta}\right)^{2}\,.$ (27)
@@
We thus find that the constants of motion $\xi$ and $\eta$ on a spherical photon orbit can be labeled by the radius $r_{p}$ of the orbit using Eqs. (32) and (33). The orbits that contribute to the critical curve on the image plane should be unstable, i.e., $d^{2}\mathcal{R}(r)/dr^{2}|_{r_{p}}>0$. By using Eqs. (32) and (33), we obtain the identity
@@
To visualize the shadow critical curve, we define the celestial coordinates $(\alpha,\beta)$ on the image plane. The coordinate $\alpha$ is the apparent perpendicular distance between the critical curve and the axis of symmetry, i.e., the spin axis. The coordinate $\beta$, on the other hand, is defined by the apparent perpendicular distance between the critical curve and the plane normal to the spin axis, with the axis piercing through the plane at the black hole. We denote the position of the observer by $(r_{0},\theta_{0})$ with $\theta_{0}$ being the inclination angle between the spin axis of the black hole and the direction to the observer. For an observer in the far exterior and asymptotically flat region $r_{0}\rightarrow\infty$, the celestial coordinates $(\alpha,\beta)$ can be expressed as Vazquez:2003zm
$\alpha$ $=\lim_{r_{0}\rightarrow\infty}\left(-r_{0}^{2}\sin\theta_{0}\frac{d\psi}{dr}\Bigg|_{r_{0},\theta_{0}}\right)\,$,
$\beta$ $=\lim_{r_{0}\rightarrow\infty}\left(r_{0}^{2}\frac{d\theta}{dr}\Bigg|_{r_{0},\theta_{0}}\right)\,.$ (35)
Using the geodesic equations Eqs. (23) and (28), one can rewrite Eq. (35) in terms of the constants of motion $(\xi,\eta)$ and the inclination angle $\theta_{0}$ in the following expressions: [^5] The expressions (36) can also be obtained using the tetrad formalism 1973ApJ…183..237C , by which one can define a generalized version of the celestial coordinates for observers at finite radius coordinates, see, e.g., Ref. Frost:2024pxa .
@@ IV.2 Shadow observables
where one integrates over the entire $\alpha$ axis on the image. One can recast this integration in terms of $r_{p}$. This integral then encompasses the entire range of unstable spherical photon orbits that contribute to the critical curve.
@@ V Results
After introducing the derivation of shadow critical curves and the associated observables, we will discuss the shadows of the black holes that satisfy the consistency of black hole thermodynamics in this section. In particular, we will focus on the black hole metric (II) with the running Newton coupling $G(r,M,a)$ parametrized by $G(z)$ with $z$ given in Eq. (17). We will consider the following sets of models: (l,p)=(1,1)(l,p)=(1,-1), $(1,0)$, $(1,4)$, and $(0,1)$. The main goal is twofold: one is to place constraints on the models using the observations of black hole images, and the other is to look for possible features in the shadow contours that are shared among these models.
@@ V.1 Corrections without mass suppression (l,p)=(1,1)(l,p)=(1,-1)
where the indices $(m,n)\in\mathbb{Z}$. In this parametrization (43), we have a set of coupling constants $(a_{i},b_{i})$ which do not depend on $M$ and are truly dimensionless. As we mentioned, the constraining power on $(a_{i},b_{i})$ will not be suppressed by the black hole mass $M$ when considering supermassive black hole images.
@@
Figure 1: The shadow critical curves (left) for $b_{1}=a_{2}=0$ and different values of $b_{2}$ and (right) for $b_{1}=b_{2}=0$ and different values of $a_{2}$.
Figure 2: Left: the critical curves for fixed $a_{2}=0$; the blue and the red curves with $(b_{1},b_{2})=(-1,0.1)$ and $(b_{1},b_{2})=(-1,-1)$, respectively. Right: the critical curves for fixed $b_{2}=0$; the blue and the red curves with $(b_{1},a_{2})=(-1,-0.1)$ and $(b_{1},a_{2})=(-1,1)$, respectively. In each panel, the shaded regions represent the results of contours scanned when taking $b_{1}\geq-1$. For $b_{1}\rightarrow\infty$, the critical curves converge to the Kerr results (black curves).
@@
As one can see from Eq. (43) that the choice of the coefficients $(a_{i},b_{i})$ may increase or decrease the Newton coupling $G(z)$, which can directly alter the size of the critical curves. In Fig. 3, we show the fractional diameter deviation $\delta$ in the parameter space $(a_{*},b_{2})$ and $(b_{1},b_{2})$ on the left and right panels, respectively, for $a_{1}=b_{1}$ and $a_{2}=0$. On the left panel, we choose $a_{1}=b_{1}=0$, while on the right panel, we fix $a_{*}=0.9$. The black and green curves indicate the constraints on $\delta$ from the upper limit of the Keck and VLTI measurements, respectively. The parameter space toward the blue region is more preferred observationally.
@@ V.2 Mass-independent scale identification $(l,p)=(1,0)$
Here, we consider the model with $(l,p)=(1,0)$, which belongs to the class of mass-independent scale identification in the FRG method. In this case, Eqs. (17) and (20) can be rewritten as
@@
Again, for generality, we assume $G(z)$ to be of the form of Padé expansion with respect to $z$,
@@
where the indices $(m,n)\in\mathbb{Z}$ and we have a corresponding set of coupling constants $(c_{i},d_{i})$. [^6] One simple mass-independent scale identification discussed in Ref. Chen:2022xjk corresponds to $c_{n}=0$ for all $n$, $d_{1}=\tilde{\omega}$, and $d_{m}=0$ for $m\neq 1$. The ansatz (47) for $G(z)$ ensures that it recovers Newton constant $G_{0}$ when $|r|\rightarrow\infty$. In fact, when incorporating Eq. (47) into the metric (II), the ADM mass is still given by $M$, and the spacetime satisfies the Solar System constraints Williams:2004qba . Also, the coupling $G(z)$ is symmetric with respect to $r=0$, although the ring on the surface $r=0$ is singular.
@@
Assuming $a_{*}=0.9$ and an edge-on inclination, i.e., $\theta_{0}=\pi/2$, we show the critical curves with different choices of $(c_{1}/G_{0}M^{2},d_{1}/G_{0}M^{2},c_{2}/G_{0}^{2}M^{4},d_{2}/G_{0}^{2}M^{4})$ in Figs. 5 and 6. We first notice that the contributions from $c_{2}$, $d_{2}$, and higher-order coupling constants are subdominant as compared with those from $c_{1}$ and $d_{1}$. One reason for this is because of the mass suppression of these higher-order corrections. Another reason is that the factor $1/(r_{*}^{2}+a_{*}^{2})$ evaluated on the spherical photon orbits is generically smaller than unity, [^7] For example, it is $1/9$ for Schwarzschild black holes and $1/2$ for extreme Kerr black holes. such that it also suppresses the higher-order corrections. Therefore, from now on, we will focus on the corrections induced by the parameters $c_{1}$ and $d_{1}$.
In addition to the hierarchy between different orders of coefficients, one can see from Figs. 5 and 6 that if the coefficients are chosen in a way such that $G(z)<G_{0}$, the critical curves shrink, acquire more distortion, and display a more D-shaped contour, e.g., the red contour in Fig. 5 where $c_{1}<0$ and the blue contour in Fig. 6 where $d_{1}>0$. Essentially, when fixing $a_{*}$, the decrease of the running Newton coupling implies a smaller event horizon, and the geometry is getting closer to the extreme configuration.
@@ V.4 Spin-independent scale identification $(l,p)=(0,1)$
Finally, we consider the spin-independent scale identification such that the Newton coupling $G(z)$ does not depend on the spin of the black hole. To satisfy the consistency of black hole thermodynamics, i.e., Eq. (15), we have $G(r,M,a)=G(Mr)$. Without loss of generality, we choose $(l,p)=(0,1)$ such that
@@
The Newton function is again parametrized using Padé expansion
@@
In Fig. 12, we show the logarithm of $\Delta C$ in the same region of the parameter space as in Fig. 11. Similar to the case of $(l,p)=(1,4)$, in some regions of the parameter space, there is only one root for $\Delta(r)=0$, and hence only one horizon exists. These regions correspond to the region above (below) the red curves in the left (right) panel. One can see that the extremality deviation $\epsilon$ can nicely capture the behavior of $\Delta C$, even within the parameter space where there is only one event horizon.
@@ V.5 Hindrance in shrinking shadow size further
To demonstrate this in our setup, which has a huge parameter space, we consider different sets of $(l,p)$, then take the spin $a_{*}$ and one of the coupling coefficients in $G(z)$ as two varying parameters, with other coupling coefficients set to zero. For a given $a_{*}$, we tune the value of the varying coupling coefficient to identify the smallest fractional diameter deviation $\delta_{\textrm{min}}$, assuming a D-shaped shadow contour. Even for a given $(l,p)$, when $a_{*}$ changes, the varying coupling coefficient that corresponds to $\delta_{\textrm{min}}$ also varies. In Fig. 13, we demonstrate the smallest fractional diameter deviation $\delta_{\textrm{min}}$ of a D-shaped shadow in terms of $a_{*}$. Each curve represents models with $\theta_{0}=\pi/2$ and different $(l,p)$ as well as the main varying coupling parameter indicated in the legend. Other coupling parameters are set to zero. In each model with a given $a_{*}$, the minimum shadow size occurs when the main varying coupling parameter is tuned such that the black hole spacetime is near the extremal configuration. When $a_{*}=1$, all models share the same $\delta_{\textrm{min}}$ with all the coupling parameters vanishing. This corresponds to the extremal configuration of the Kerr black hole. On the other hand, when $a_{*}=0$, the model of (l,p)=(1,1)(l,p)=(1,-1) with varying $a_{2}$ (magenta solid curve), the model of $(l,p)=(1,0)$ with varying $c_{1}$ (blue solid curve), and the model of $(l,p)=(0,1)$ with varying $g_{2}$ (black solid curve) degenerate because they both have $G(z)/G_{0}-1\propto 1/r^{2}$. Similarly, the model of $(l,p)=(1,0)$ with varying $d_{1}$ (blue dashed curve), the model of $(l,p)=(0,1)$ with varying $h_{2}$ (black dashed curve), and the model of (l,p)=(1,1)(l,p)=(1,-1) with varying $b_{2}$ (magenta dashed curve) degenerate because they all have $G_{0}/G(z)-1\propto 1/r^{2}$. The shaded regions surrounding the curves represent the results scanned for varied $\theta_{0}$.
@@ VI Spinning black holes beyond Kerr bound
In Secs. V.3 and V.4 in which we considered the models with $(l,p)=(1,4)$ and $(l,p)=(0,1)$, respectively, we pointed out that there are regions of parameter space where $\Delta(r)=0$ only has one root, hence only one horizon exists. One natural question is as follows: in these cases, will the event horizon still exist if one keeps increasing the spin parameter $a_{*}$ and goes beyond the usual Kerr bound, i.e., $a_{*}\leq 1$?
@@
It turns out that it is possible to go beyond the Kerr bound, but the shadow images may develop irregular structures. More explicitly, we find that when $a_{*}$ is sufficiently large, a branch of stable spherical photon orbits may appear outside the event horizon. This branch of stable spherical photon orbits connects two distinct branches of unstable photon orbits. In this case, the shadow critical curve develops a cuspy structure, as shown in Fig. 14, i.e., the contours with $a_{*}=3$ and $a_{*}=4$. This can be understood from Fig. 15, where we show the constants of motion $\xi(r_{p})$ and $\eta(r_{p})$ given in Eqs. (32) and (33), respectively. As we discussed in Sec. IV, these constants of motion are functions of the radii $r_{p}$ of different spherical photon orbits. Note that in this figure, the horizontal axis is expressed by the rescaled radius $r_{p*}\equiv r_{p}/G_{0}M$. On the left panel of Fig. 15, we choose the parameters that correspond to the darker green contour in Fig. 14, i.e., $a_{*}=2$. In this case, $\xi(r_{p})$ is a monotonically decreasing function in $r_{p}$, and the orbits that contribute to the contour are all unstable. The shadow critical curve is a smooth and closed contour. On the other hand, on the right panel of Fig. 15, in which the spin $a_{*}=4$ is sufficiently large, $\xi(r_{p})$ has one local minimum and one local maximum, which are connected by a branch of stable spherical photon orbits (the dashed green segment; see also Eq. (34)). This stable branch of orbits connects to two separated branches of unstable orbits on its two ends. There are two distinct unstable orbits, each belonging to the two sides of unstable branches, that contribute to the same point $(\alpha,\beta)$ on the image plane, i.e., the cusps on the contours. Such a discontinuity gives rise to the cuspy structure in the shadow critical curve, and it is a manifestation of the rich spacetime structure in this model, albeit the Liouville integrability of the geodesic dynamics ensured by the Carter constant $\mathcal{K}$ (see Eqs. (28) and (29)). Black hole models with similar cuspy structures in the shadow critical curves can also be found in Refs. Cunha:2017eoe ; Wang:2017hjl . [^10] Cuspy structure in shadow boundary has also been reported in the models in which two distinct branches of unstable photon orbits transit discontinuously, without resorting to stable spherical photon orbits Qian:2021qow . The model considered in Ref. Wang:2017hjl also has Liouville integrable geodesics. The metric line element of the model considered in Ref. Wang:2017hjl can be expressed in the form of Eq. (II) with $G(r)=G_{0}+\bar{\eta}/2Mr^{2}$ and $\bar{\eta}$ a non-Kerr parameter. This Newton coupling satisfies thermodynamic consistency only when $\bar{\eta}\propto 1/M$. If this choice is made, this falls into the class of solutions discussed in Sec. V.4.
@@
Before closing this section, we would like to comment on the physical implications of this type of cuspy shadows and the corresponding black hole spacetimes. First of all, the shadow contours with such a ``circular sector" shape are very different from the usual D-shaped shadows, and can be easily distinguished through observations. Furthermore, from theoretical points of view, the existence of stable spherical photon orbits implies the possibility of accumulating photons (or gravitons) around the black hole for a long time. Such an accumulation may hint toward some sorts of instability of the entire geometry. The relevant questions are whether such systems are truly unstable and, if so, how long the instability timescale is. However, the answer to these questions requires a thorough numerical analysis that takes into account the nonlinear effects of the underlying gravitational theory, as well as the backreactions induced by the accumulating particles. Such an analysis is beyond the scope of this paper.
@@
Figure 15: Left: for the black holes whose shadow critical curves do not have cuspy structures, the constant of motion $\xi(r_{p})$ is a monotonic function of $r_{p}$. All the spherical photon orbits that contribute to the shadow critical curve are unstable. Right: the cuspy structure on the image plane in Fig. 14 is due to the existence of the local maximum and the minimum in $\xi(r_{p})$. In particular, the spherical photon orbits with radii between the two local extrema, i.e., the green dashed line segment, are stable orbits with $d^{2}\mathcal{R}(r)/dr^{2}|_{r_{p}}<0$ (see Eq. (34)). The impact parameters of the orbits in this line segment are represented by the dashed segments in Fig. 14. Note that for $\theta_{0}=\pi/2$, the spherical photon orbits that contribute to the critical curve are those with radii bounded by the two roots of $\eta(r_{p})=0$. The left edge of each panel represents the location of the event horizon, which is very close to the smaller root of $\eta(r_{p})=0$.